• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 16:01:58 on October13,2005

    Entry number 156745

    keyword=Day Shift Summary



    Shift Summary: Day Shift October 13 2005


    HAPPEX-II hydrogen

    Shift Leader/Target:   D.S. Armstrong
    Target:   R. Gilman
    3rd:   J. LeRose



       At start of shift: Kent is here as well; just at start of shift we get 5 µA of beam, back to nominal beam positions.
    We have target full, but in empty position. Take a HAPPEX DAQ run to
    look at phase advance and dithering. We are running at the moment
    with the Hall B laser, as the Hall A laser will not lock, and EGG needs to do an access to fix that today.

    8:30 We check the phase advance with 5 µA beam.
    Looks good in x,y - Kent pronounces it fine after pumping up the amplitude in one coil. Now MCC takes beam to allow Matt Poelker to fix the Hall A laser.

    Pedestal/cosmic runs underway while we wait.

    10:00 Elog says Matt has fixed the Hall A laser, and it is now
    locked. Now they are going to the "energy conservation" test plan (in
    which the higher-pass arc magnet power supplies are turned off). Matt will also be resetting the Wein filter/precession solenoid magnet to zero-out the vertical polarization.

    16:00 Still waiting for beam... Hand off to next shift...



    HAPPEX Runs Taken

    Run Pairs Comment

    5588 45 K beam study - phase advance check
    5589 100 K Cosmics/Pedestals
    5590 100 K Cosmics/Pedestals
    5591 100 K Cosmics/Pedestals
    5592 100 K Cosmics/Pedestals
    5593 100 K Cosmics/Pedestals
    5594 100 K Cosmics/Pedestals
    5595 100 K Cosmics/Pedestals
    5596 100 K Cosmics/Pedestals
    5597 100 K Cosmics/Pedestals